OpenStudy (punderson):

2. The tables below show the values of y corresponding to different values of x: Table A x 3 3 2 y 1 0 0 Table B x 3 5 5 y -2 2 -2 Which statement is true for the tables? (1 point) Both Table A and Table B represent functions. Both Table A and Table B do not represent functions. Table A does not represent a function, but Table B represents a function. Table A represents a function, but Table B does not represent a function. @umerlodhi

OpenStudy (bmk614):

Both table A and Table B do not represent functions because when using the horizontal line test, y can not equal more than one x.

OpenStudy (punderson):

3. Which of the following equations best represents the graph of f(p)? (1 point) y = f(p) f(y) = p f(x) = p x = f(p)

umerlodhi (umerlodhi):

i suck at math but ill try

OpenStudy (bmk614):

y=f(p)

OpenStudy (punderson):

Thats wrong @bmk614